a coin is placed on a rotating disc whose angular velocity is 2 pi rad/s. If the coefficient of static friction between the coin and the disc is 0.6, what is the maximum distance from the center of the disc that the coin can be placed on without it slipping off?
Answer #1

Friction provides required centripetal force.

μmg=mrΩ2

r=μg÷Ω^2=0.6×9.8÷(2×3.14)^2=0.149 m